# ConKeD: Multiview contrastive descriptor learning for keypoint-based retinal image registration

### v1

@@ 2.1.3 Keypoint matching and transformation computation
After the keypoint extraction, the keypoint descriptors are computed by the descriptor network trained using ConKeD, which produces a dense block of descriptors for each image. Using the keypoints’ locations, the descriptors for each keypoint are selected. Next, these descriptors are matched among the images in each registration pair. It should be noted that, for a descriptor to match another one, this match should be bidirectional. That is, a given descriptor A matches with another descriptor B, if both A is the closest to B and B is the closest to A. As in the training stage, the distance between descriptors is measured using cosine similarity. Additionally, as the keypoint detection network can accurately classify the keypoints in crossovers of bifurcations, we leverage this capability to ease and speed up the matching. In particular, we only match the descriptors of points within each class, that is, crossovers with crossovers and bifurcations with bifurcations. This drastically reduces the amount of required computations, which was already low due to the high specificity of these keypoints.
After the descriptor matching step, the paired keypoints are used to estimate the geometric transformation using RANSAC Fischler and Bolles (1981). This is a well known algorithm that is commonly used to estimate the parameters of mathematical models when there may be outliers in the data. In this case, the model is the transformation matrix between the images and the data are the matched keypoints. We employ a projective (or homographical) transformation, as previous works in the state-of-the-art Rivas-Villar et al. ; Liu et al. (2022).
@@ 2.2.1 Dataset
For the training phase we employ the public DRIVE dataset. The images were obtained from a diabetic retinopathy screening program using a Canon CR5 non-mydriatic 3CCD camera with a FOV of 45 degrees. Each image has a resolution of $584\times 565$ pixels. This dataset of 40 images is equally divided in the training and test sets, each of 20 images. This dataset has a ground truth of vessel crossovers and bifurcations that we use to train our keypoint detector Abbasi-Sureshjani et al. (2016). The original binary labeling from DRIVE is converted to heatmaps, as explained in Álvaro S. Hervella et al. (2020a).
@@ 2.2.2 Keypoint detection
We train the keypoint detector network as in Álvaro S. Hervella et al. (2020a). That is, we use the training set of the DRIVE dataset (i.e. 20 images), while reserving 25% of the images for the validation set. The remaining 20 images compose the test set. For this task, we use the U-Net Ronneberger et al. (2015) network architecture as in Álvaro S. Hervella et al. (2020a). To train the network, we use Adam Kingma and Ba (2015) with learning rate decay. The network was trained from scratch. The learning rate is originally set to $41e-$ 4 with a patience of 2500 batches before reducing it. Each batch contains a single image, and each time the learning rate is reduced, it is by a factor of 0.1. The training stops after the learning rate reaches $71e-$ 7. We use spatial augmentation consisting of random affine transformations with the following parameters: random rotations of ±$\pm 90^{\circ}$, random scaling between $0.9-1.1\times imageSize$ and random shearing of ±$\pm 20^{\circ}$. Finally, color augmentation is also used randomly changing image components in the HSV color space Liskowski and Krawiec (2016).
The value for the intensity threshold was selected by testing multiple values and finding the one that provided the best F1-Score in the test set of the DRIVE dataset. In this case, 0.35 was found to be the optimal value Álvaro S. Hervella et al. (2020a).
@@ 2.2.3 Keypoint description
For the training, we use the whole DRIVE dataset (i.e. all the 40 images) in combination with the keypoint output produced by the detection network. Therefore, this network is trained using the DRIVE dataset resolution. As network architecture, we use the modified L2-net Tian et al. (2017) proposed in R2D2 Revaud et al. (2019). This network produces a descriptor per input pixel without the need to upscale the descriptor block. Moreover, it has demonstrated to produce accurate descriptors in other works Revaud et al. (2019). We train the network from scratch for 15000 epochs using Adam Kingma and Ba (2015) with a constant learning rate of $41e-$ 4. For both losses, SupCon Loss and MP-InfoNCE, we set the temperature parameter $\tau=0.1$. As augmentations we use random affine transformations with rotations of ±$\pm 60^{\circ}$, translations of $0.25\times imageSize$ in each axis, scaling between $0.75-1.25\times imageSize$ and shearing of ±$\pm 30^{\circ}$. Moreover, we also include color augmentation in the form of random changes in the HSV colorspace Álvaro S. Hervella et al. (2020b) as well as random Gaussian Noise with a mean of 0 and a standard deviation of 0.05. The geometric and color augmentations are applied to every augmented image while the random Gaussian noise is applied with a probability of 0.25. During training, the detection of keypoints is performed only on the original images. Therefore, the geometric augmentations are applied to both images and keypoints. Finally, regarding the multiviewed batch sizes, we perform experiments with 3 different sizes. Using a single original image, we try the multiviewed batch sizes 1+1 (i.e., 1 original image and 1 augmentation), 1+9, and 1+19. Moreover, we also train a descriptor network using the triplet loss, to use as baseline. In order for the comparison to be completely fair we use cosine similarity as the distance metric. Moreover, we use a margin of value 0.05, which we empirically found to be the most adequate. The training is done using the exact same settings described above except that, in this case, we use a learning rate of $51e-$ 5 and train the network during 75000 epochs.
@@ 2.2.4 Keypoint matching
After the description step, we match the descriptors for both images composing a registration pair. The resulting keypoint matches are upscaled from the detection resolution (DRIVE) to the test resolution (FIRE). Once the keypoints are upscaled, they are used in combination with RANSAC Fischler and Bolles (1981) to create a projective transformation that aligns the image pair. Therefore, the test images are aligned using their full size which is required to compare among methods.
@@ 2.3 Evaluation methodology
To validate our approach, first we use the evaluation metric proposed by the authors of the FIRE dataset Hernandez-Matas et al. (2017). This metric, called Registration Score, measures the average Euclidean distance between the control points of each image pair after the registration process. Then, depending if the error of an image pair is below a threshold or not, that pair can be deemed successful or unsuccessful. By plotting the ratio of successful registrations on the Y axis and progressively increasing the error threshold along the X axis, we can create a 2D graph and compute its Area Under Curve (AUC). This AUC, specifically computed between 0 and 25 pixels of error threshold, is the registration score that we use as metric to compare methods. The AUC as well as the 2D graph allow for an easy and robust comparison among the different methods evaluated on the FIRE dataset. In that regard, it should be noted that Registration Score has been widely used as a robust metric to evaluate image registration in many state of the art works Rivas-Villar et al. (2022); Hernandez-Matas et al. (2020); Motta et al. (2019); Liu et al. (2022). We calculate this metric for each category separately as well as for the whole dataset as proposed in FIRE Hernandez-Matas et al. (2017). Moreover, in order to compare our results to Liu et al. (2022) we also compute the average (Avg.) of the registration score among the three categories. However, as this does not take into account the varying number of images per category, we also propose to compute a weighted average (W. Avg.) among the three categories, using the number of images per category as the weighting factor. This metric should provide a more representative picture of the performance through the whole dataset.
@@
The Registration Score only considers the error produced using all the detected keypoints. Thus, this metric may encourage brute-forcing both detection and description to improve the results, without considering the efficiency or computation time. Given an accurate enough description, simply detecting more points (i.e, higher keypoint frequency) improves the chance that each keypoint matches with another one in its approximate corresponding space, even if the detector is pseudo-random. However, this entails a much higher computational cost. Performing the registration with a smaller number of keypoints (i.e. more spaced out) puts more emphasis on both steps, as the detection needs to be repeatable enough so that the keypoints are in approximately the same locations in both image of a pair. Thus, using more keypoints increases the chance of accurate registration by pure brute force, increasing computational complexity, while using less, more relevant, keypoints is beneficial due to the lower computational cost.
@@
In order to perform a more complete evaluation, we propose a new metric based on the registration score that takes into account the number of matching keypoints. By limiting the amount of paired keypoints (i.e. after descriptor matching) used in RANSAC we can better evaluate the quality of the detected keypoints and their descriptors. For this metric we propose to use the top-N keypoints, as ordered by the distance of their descriptors. That is, we take the N keypoints whose descriptors are closest, i.e. which offer a higher quality match. By increasing the value of N (i.e. the number of keypoint pairs) and computing the registration score for each one of the values, we can plot another curve and calculate its AUC. This way, we can evaluate the performance of the network across different number of keypoints. It should be noted that we can not have less keypoints than the required by the transformation model, therefore $N>4$. Moreover, as we match the keypoints independently per class (i.e. crossovers and bifurcations), we compute the top-N points for each class. In particular, we start the evaluation with the top-3 keypoints per class, which is a total of 6 keypoints (i.e. the best 3 crossovers and the best 3 bifurcations). Next we increase the number to top-4, which is 8 points in total and so on, up to top-25 (with a total of 50 keypoints). In this evaluation we set the RANSAC budget according to the amount of keypoints such that all RANSAC possibilities are computed. It should be noted that, in some cases, the value of N might be bigger than the number of available or matched keypoints for each class (either crossovers or bifurcations). In this case no more points are added to that specific class. We name this metric Variable Top Keypoint Registration Score (VTKRS). Similarly to the registration score, we also compute the VTKRS per category and overall for the whole dataset.
@@ 3 Results and discussion
This section is structured as follows: first, in subsection 3.1 we evaluate our method. This includes a comparison with a triplet baseline, as well as analysing the effect of including multiple positives as well as the effect of increasing the number of positives, by increasing the multiviewed batch size. Next, in subsection 3.2 we compare both of the proposed loss functions using the best multiviewed batch size. Finally, in subsection 3.3, we compare our approach with the different state of the art methods.
@@ 3.1 Evaluation of the proposed approach
In order to evaluate the performance of ConKeD, we compare our approach to the Triplet Loss, used in state-of-the art approaches Liu et al. (2022). Additionally, we compare three different variants of our proposal, the first one using a batch of 1+1 views such that our method is multi-negative and single-positive and the other two using batches of 1+9 and 1+19 views each such that our method is both multi-negative and multi-positive. Therefore, we are comparing a single-negative single-positive (triplet) against a multi-negative single-positive (1+1 views) to the novel multi-positive multi-negative methods (1+9, 1+19). For this comparison we use SupCon loss Khosla et al. (2020). Table 1 shows the results for this experiment.
Firstly, the results show that our proposal clearly outperforms the triplet loss approach, regardless of the multiviewed batch size. Moreover, we have empirically found that training with multi-positive multi-negative losses (either SupCon loss or MP-InfoNCE) drastically reduces the training time, requiring five times less iterations to converge than the triplet loss. In terms of categories in the FIRE dataset, the difference between methods is less notorious in the category S, which is more forgiving with inaccurate descriptor matches due to the high amount of overlapping. On the other hand, the difference in categories A and P is notable due to the disease progression and low overlapping. These characteristics make accurate descriptor matching more relevant, due to the lower amount of available keypoints. In the category P, there are less keypoints in the overlapping zone and, therefore, less usable keypoints in the registration. Thus, it is more important to correctly match the few usable keypoints, in order to be able to produce an accurate alignment. In category A, due to the progress of pathologies, the amount of keypoints is also relevant, as some may be vanished or obstructed. However, given the higher amount of overlapping, the number of keypoints is less limiting, putting more weight on the robust description that allows to recognize keypoints under different conditions (i.e., pathological progression). Therefore, to evaluate the descriptors, category A is the most relevant one. It should be noted that, as the detected keypoints are exactly the same in all the approaches, the differences are purely due to the improved descriptors.
@@
Regarding the number of views in a multi-positive setting, (i.e., 1+9 and 1+19) the results are equivalent with both metrics. Moreover, given the extended training in our experimentation, i.e., 15000 epochs, we can ensure that the difference in performance between 1+1 and other two approaches is not due to the different number of seen image views. All three approaches are trained well beyond their convergence point, such that, there are no more meaningful improvements possible. This serves to isolate the improvements of the multi-positive approach from the potential improvements of using a higher number of views per batch and thus allowing the network to see more images per iteration. Therefore, the difference between adding multiple positives (1+1 vs the rest) and just increasing the number of views in the batch (1+9 vs 1+19) is clear. Given that the results of batches 1+9 and 1+19 are equivalent, we choose 1+9 as the reference batch to compare losses and models in the following sections.
@@ 3.2 Comparison of loss functions
In this section, we compare SupCon Loss against MP-InfoNCE. The experiments are performed using the optimal multiviewed batch size found in the previous experiments (i.e. 1+9). The results are detailed in Tables 3 and 4 as well as in Fig. 4.
In terms of registration score over the FIRE dataset, detailed in Table 3, the results for both losses are very similar. The biggest differences appear in Category A ($+1.1\%$ for SupCon) and Category P ($+1.2\%$ for MP-InfoNCE). However, these differences are small due to the limited nature of the registration score evaluation over FIRE. The results in terms of VTKRS can be seen in Table 4 and in Fig. 4. The final numeric differences in the AUC values for both losses are small, like in the previous evaluation. However looking at the curves, the differences are much more evident. In particular, for category A, we can see that MP-InfoNCE provides results that are more accurate with a low number of keypoints. In category P both losses perform similarly, MP-InfoNCE obtains better results with less points while SupCon Loss improves more given more points. However, in the end, the two approaches become virtually even when many or all of the keypoints are used.
The difference in performance in the Category A (the only one where the difference is notable) is specially relevant as this is the category with the most clinical relevance. RIR is commonly used for disease monitoring and study follow-up, therefore, any automated method should desirably be robust to pathological lesions. Additionally, Category A can be viewed as the hardest category to align due to the number and variety of lesions and diseases that severely alter the morphology of the ocular fundus. While Category P obtains the lowest scores, this is due to the lower overlapping and thus the lower amount of common keypoints on both images of the registration pair. However, in category A the detected keypoints can disappear or be altered by pathology progression. Therefore, improving the results in this category is desirable and even more with a low number of keypoints as this makes the network more robust to pathology progression which can completely occlude the keypoints. Given this benefit of MP-InfoNCE over SupCon Loss we will use the MP-InfoNCE as the method to compare our approach against the state of the art. Nevertheless, while there are differences between MP-InfoNCE and SupCon Loss, both loss terms produce satisfactory performance. Therefore, we can conclude that our multi-negative multi-positive approach is robust to the specific loss term used.

### v2

@@ 2.1.3 Keypoint matching and transformation computation
After the keypoint extraction, the keypoint descriptors are computed by the descriptor network trained using ConKeD, which produces a dense block of descriptors for each image. Using the keypoints’ locations, the descriptors for each keypoint are selected. Next, these descriptors are matched among the images in each registration pair. It should be noted that, for a descriptor to match another one, this match should be bidirectional. That is, a given descriptor A matches with another descriptor B, if both A is the closest to B and B is the closest to A. As in the training stage, the distance between descriptors is measured using cosine similarity. Additionally, as the keypoint detection network can accurately classify the keypoints in crossovers of bifurcations, we leverage this capability to ease and speed up the matching. In particular, we only match the descriptors of points within each class, that is, crossovers with crossovers and bifurcations with bifurcations. This drastically reduces the number of required computations, which was already low due to the high specificity of these keypoints.
After the descriptor matching step, the paired keypoints are used to estimate the geometric transformation using RANSAC [22]. This is a well-known algorithm that is commonly used to estimate the parameters of mathematical models when there may be outliers in the data. In this case, the model is the transformation matrix between the images and the data are the matched keypoints. We employ a projective (or homographical) transformation, as previous works in the state-of-the-art [8, 24].
@@ 2.2.1 Dataset
For the training phase we employ the public DRIVE dataset. The images were obtained from a diabetic retinopathy screening program using a Canon CR5 non-mydriatic 3CCD camera with a FOV of 45 degrees. Each image has a resolution of $584\times 565$ pixels. This dataset of 40 images is equally divided in the training and test sets, each of 20 images. This dataset has a ground truth of vessel crossovers and bifurcations that we use to train our keypoint detector [35]. The original binary labeling from DRIVE is converted to heatmaps, as explained in [33].
@@ 2.2.2 Keypoint detection
We train the keypoint detector network as in [33]. That is, we use the training set of the DRIVE dataset (i.e. 20 images), while reserving 25% of the images for the validation set. The remaining 20 images compose the test set. For this task, we use the U-Net [36] network architecture as in [33]. To train the network, we use Adam [37] with learning rate decay. The network was trained from scratch. The learning rate is originally set to $41e-$ 4 with a patience of 2500 batches before reducing it. Each batch contains a single image, and each time the learning rate is reduced, it is by a factor of 0.1. The training stops after the learning rate reaches $71e-$ 7. We use spatial augmentation consisting of random affine transformations with the following parameters: random rotations of ±$\pm 90^{\circ}$, random scaling between $0.9-1.1\times imageSize$ and random shearing of ±$\pm 20^{\circ}$. Finally, color augmentation is also used randomly changing image components in the HSV color space [38].
The value for the intensity threshold was selected by testing multiple values and finding the one that provided the best F1-Score in the test set of the DRIVE dataset. In this case, 0.35 was found to be the optimal value [33].
@@ 2.2.3 Keypoint description
For the training, we use the whole DRIVE dataset (i.e. all the 40 images) in combination with the keypoint output produced by the detection network. Therefore, this network is trained using the DRIVE dataset resolution. As network architecture, we use the modified L2-net [39] proposed in R2D2 [29]. This network produces a descriptor per input pixel without the need to upscale the descriptor block. Moreover, it has demonstrated to produce accurate descriptors in other works [29]. We train the network from scratch for 15000 epochs using Adam [37] with a constant learning rate of $41e-$ 4. For both losses, SupCon Loss and MP-InfoNCE, we set the temperature parameter $\tau=0.1$. As augmentations we use random affine transformations with rotations of ±$\pm 60^{\circ}$, translations of $0.25\times imageSize$ in each axis, scaling between $0.75-1.25\times imageSize$ and shearing of ±$\pm 30^{\circ}$. Moreover, we also include color augmentation in the form of random changes in the HSV color space [40] as well as random Gaussian Noise with a mean of 0 and a standard deviation of 0.05. The geometric and color augmentations are applied to every augmented image while the random Gaussian noise is applied with a probability of 0.25. During training, the detection of keypoints is performed only on the original images. Therefore, the geometric augmentations are applied to both images and keypoints. Finally, regarding the multiviewed batch sizes, we perform experiments with 3 different sizes. Using a single original image, we try the multiviewed batch sizes 1+1 (i.e. 1 original image and 1 augmentation), 1+9, and 1+19. Moreover, we also train a descriptor network using the triplet loss, to use as baseline. In order for the comparison to be completely fair we use cosine similarity as the distance metric. Moreover, we use a margin of value 0.05, which we empirically found to be the most adequate. The training is done using the exact same settings described above except that, in this case, we use a learning rate of $51e-$ 5 and train the network during 75000 epochs.
@@ 2.2.4 Keypoint matching
After the description step, we match the descriptors for both images composing a registration pair. The resulting keypoint matches are upscaled from the detection resolution (DRIVE) to the test resolution (FIRE). Once the keypoints are upscaled, they are used in combination with RANSAC [22] to create a projective transformation that aligns the image pair. Therefore, the test images are aligned using their full size which is required to compare among methods.
@@ 2.3 Evaluation methodology
To validate our approach, first we use the evaluation metric proposed by the authors of the FIRE dataset [41]. This metric, called Registration Score, measures the average Euclidean distance between the control points of each image pair after the registration process. Depending on whether the error of an image pair falls below a threshold, the pair is classified as either successful or unsuccessful. By plotting the ratio of successful registrations on the Y axis and progressively increasing the error threshold along the X axis (defined between 1 and 25 pixels for FIRE), we can create a 2D graph and compute its Area Under Curve (AUC), allowing for quick and easy comparison among methods [23, 20, 21, 24].
We calculate this metric for each category separately as well as for the whole dataset as proposed in FIRE [41]. Moreover, in order to compare our results to [24] we also compute the average (Avg.) of the registration score among the three categories. We also propose to compute a weighted average (W. Avg.) among the three categories, using the number of images per category as the weighting factor. This metric should provide a more representative picture of the performance through the whole dataset.
@@
The Registration Score only considers the error produced using all the detected keypoints. Thus, this metric may encourage brute-forcing both detection and description to improve the results, without considering the efficiency or computation time. Given an accurate enough description, simply detecting more points (i.e. higher keypoint frequency) improves the chance that each keypoint matches with another one in its approximate corresponding space, even if the detector is pseudo-random. However, this entails a much higher computational cost. Performing the registration with a smaller number of keypoints (i.e. more spaced out) puts more emphasis on both steps, as the detection needs to be repeatable enough so that the keypoints are in approximately the same locations in both image of a pair.
@@
In order to perform a more complete evaluation, we propose a new metric based on the registration score that takes into account the number of matching keypoints. By limiting the amount of paired keypoints (i.e. after descriptor matching) used in RANSAC we can better evaluate the quality of the detected keypoints and their descriptors. For this metric we propose to use the top-N keypoints, as ordered by the distance of their descriptors. That is, we take the N keypoints whose descriptors are closest, i.e. which offer a higher quality match. By increasing the value of N (i.e. the number of keypoint pairs) and computing the registration score for each one of the values, we can plot another curve and calculate its AUC. This way, we can evaluate the performance of the network across different number of keypoints. It should be noted that we cannot have less keypoints than the required by the transformation model, therefore $N>4$. Moreover, as we match the keypoints independently per class (i.e. crossovers and bifurcations), we compute the top-N points for each class. In particular, we start the evaluation with the top-3 keypoints per class, which is a total of 6 keypoints (i.e. the best 3 crossovers and the best 3 bifurcations). Next, we increase the number to top-4, which is 8 points in total and so on, up to top-25 (with a total of 50 keypoints). In this evaluation we set the RANSAC budget according to the amount of keypoints such that all RANSAC possibilities are computed. It should be noted that, in some cases, the value of N might be bigger than the number of available or matched keypoints for each class (either crossovers or bifurcations). In this case no more points are added to that specific class. We name this metric Variable Top Keypoint Registration Score (VTKRS). Similarly to the registration score, we also compute the VTKRS per category and overall for the whole dataset.
@@ 3 Results and discussion
This section is structured as follows: first, in subsection 3.1 we evaluate our method. This includes a comparison with a triplet baseline, as well as analyzing the effect of including multiple positives as well as the effect of increasing the number of positives, by increasing the multiviewed batch size. Next, in subsection 3.2 we compare both of the proposed loss functions using the best multiviewed batch size. Finally, in subsection 3.3, we compare our approach with the different state of the art methods.
@@ 3.1 Evaluation of the proposed approach
In order to evaluate the performance of ConKeD, we compare our approach to the Triplet Loss, used in state-of-the art approaches [24]. Additionally, we compare three different variants of our proposal, the first one using a batch of 1+1 views such that our method is multi-negative and single-positive and the other two using batches of 1+9 and 1+19 views each such that our method is both multi-negative and multi-positive. Therefore, we are comparing a single-negative single-positive (triplet) against a multi-negative single-positive (1+1 views) to the novel multi-positive multi-negative methods (1+9, 1+19). For this comparison we use SupCon Loss [30]. Table 1 shows the results for this experiment.
Firstly, the results show that our proposal clearly outperforms the triplet loss approach, regardless of the multiviewed batch size. Moreover, we have empirically found that training with multi-positive multi-negative losses (either SupCon Loss or MP-InfoNCE) drastically reduces the training time, requiring five times less iterations to converge than the triplet loss. In terms of categories in the FIRE dataset, the difference between methods is less notorious in the category S, which is more forgiving with inaccurate descriptor matches due to the high amount of overlapping. On the other hand, the difference in categories A and P is notable due to the disease progression and low overlapping. These characteristics make accurate descriptor matching more relevant, due to the lower amount of available keypoints. In the category P, there are less keypoints in the overlapping zone and, therefore, less usable keypoints in the registration. Thus, it is more important to correctly match the few usable keypoints, in order to be able to produce an accurate alignment. In category A, due to the progression of pathologies, the amount of keypoints is also relevant, as some may be vanished or obstructed. However, given the higher amount of overlapping, the number of keypoints is less limiting, putting more weight on the robust description that allows to recognize keypoints under different conditions (i.e. pathological progression). Therefore, to evaluate the descriptors, category A is the most relevant one. It should be noted that, as the detected keypoints are exactly the same in all the approaches, the differences are purely due to the improved descriptors.
@@
Regarding the number of views in a multi-positive setting, (i.e. 1+9 and 1+19) the results are equivalent with both metrics. Moreover, given the extended training in our experimentation, i.e. 15000 epochs, we can ensure that the difference in performance between 1+1 and other two approaches is not due to the different number of seen image views. All three approaches are trained well beyond their convergence point, such that, there are no more meaningful improvements possible. This serves to isolate the improvements of the multi-positive approach from the potential improvements of using a higher number of views per batch and thus allowing the network to see more images per iteration. Therefore, the difference between adding multiple positives (1+1 vs the rest) and just increasing the number of views in the batch (1+9 vs 1+19) is clear. Given that the results of batches 1+9 and 1+19 are equivalent, we choose 1+9 as the reference batch to compare losses and models in the following sections.
@@ 3.2 Comparison of loss functions
In this section, we compare SupCon Loss against MP-InfoNCE. The experiments are performed using the optimal multiviewed batch size found in the previous experiments (i.e. 1+9). The results are detailed in Tables 3 and 4 as well as in Fig. 5. Representative examples of registered images from the FIRE dataset using our both losses can be seen in the top rows of Figure 6.
In terms of registration score over the FIRE dataset, detailed in Table 3, the results for both losses are very similar. The biggest differences appear in Category A ($+1.1\%$ for SupCon) and Category P ($+1.2\%$ for MP-InfoNCE). However, these differences are small due to the limited nature of the registration score evaluation over FIRE. The results in terms of VTKRS can be seen in Table 4 and in Fig. 5. The final numeric differences in the AUC values for both losses are small, like in the previous evaluation. However, looking at the curves, the differences are much more evident. In particular, for category A, we can see that MP-InfoNCE provides results that are more accurate with a low number of keypoints. In category P both losses perform similarly, MP-InfoNCE obtains better results with less points while SupCon Loss improves more given more points. However, in the end, the two approaches become virtually even when many or all of the keypoints are used.
The difference in performance in the Category A (the only one where the difference is notable) is especially relevant as this is the category with the most clinical relevance. RIR is commonly used for disease monitoring and study follow-up, therefore, any automated method should desirably be robust to pathological lesions. Additionally, Category A can be viewed as the hardest category to align due to the number and variety of lesions and diseases that severely alter the morphology of the ocular fundus. While Category P obtains the lowest scores, this is due to the lower overlapping and thus the lower amount of common keypoints on both images of the registration pair. However, in category A the detected keypoints can disappear or be altered by pathology progression. Therefore, improving the results in this category is desirable and even more with a low number of keypoints as this makes the network more robust to pathology progression which can completely occlude the keypoints. Given this benefit of MP-InfoNCE over SupCon Loss we will use the MP-InfoNCE as the method to compare our approach against the state of the art. Nevertheless, while there are differences between MP-InfoNCE and SupCon Loss, both loss terms produce satisfactory performance. Therefore, we can conclude that our multi-negative multi-positive approach is robust to the specific loss term used.